# Sample Compression for Self-Certified Continual Learning

### v3

@@ Appendix A Compression and Reconstruction Functions
19:  return $\theta_{T},\{S^{\operatorname{\mathbf{i}}}_{t}\}_{t=1}^{T},\{S^{\operatorname{\mathbf{j}}}_{t}\}_{t=1}^{T},\{\mu_{1,t}\}_{t=1}^{T},\mu_{2}$. {Learned parameters, compression sets and messages sets}
A.1 Compression Function
Two challenges prevent CoP2L (as presented by Algorithm 3) from being used as its own compression and reconstruction function. In this subsection, we expose each of these challenges while explaining how we the proposed compression algorithm (Algorithm 4) allows to circumvent them.
The first challenge comes from the sampling of the buffer $B^{\star}$ (Lines 6 to 8 of Algorithm 3). When training using CoP2L, datapoints from the buffer can be chosen to be part of the compression set. At the next task, these datapoints may still be available to train the model (if they are not excluded from the buffer by the sampling step). Therefore, the reconstruction function needs to know when to remove the datapoint. The message $\mu_{1}$ returned by Algorithm 4 thus provides the task index after which the datapoint was removed during CoP2L execution. As the reconstruction function only requires a message for datapoints who were removed from the buffer after being added to the compression set $S^{\operatorname{\mathbf{i}}}$, a second compression set $S^{\operatorname{\mathbf{j}}}$ is dedicated to datapoints that necessitate a message. The message $\mu_{1}$ given is chosen among $\{2,\ldots,T\}$ for each datapoint belonging to $S^{\operatorname{\mathbf{j}}}$.
The second challenge is the use of the bound as stopping criterion in mP2L. Recall that the reconstruction function does not have access to the whole dataset. Thus, it cannot recover the bound value computed during the initial training phase (see Line 11 of Algorithm 2) and use it as a stopping criterion. To address this inconvenience, the number of iterations to perform is provided to mP2L as a message $\mu_{2}=(\mu_{2}^{1},\ldots,\mu_{2}^{T})$. That is, for each task $t$, with a buffer $\mathcal{B}$, the message component $\mu_{2}^{t}$ is chosen in $\{1,\ldots,n_{t}+|\mathcal{B}|\}$, giving the number of mP2L’s iterations to perform.
@@ A.1 Compression Function
$\mathscr{M}(\operatorname{\mathbf{j}},T)\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\textstyle\bigtimes_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (7)
A.2 Reconstruction Function
Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function (Algorithm 4) and for CoP2L (Algorithm 3), the composition of the compression and the reconstruction function must output the exact same predictor as CoP2L. However, it only has access to the compression sets and the messages. The first challenge of the compression function is addressed using $\mu_{1}$ from line 6 to 11 of Algorithm 5. For each datasets, we verify if a datapoint was previously excluded from the buffer by the sampling function. If so, we remove it from the buffer. The second challenge is addressed in line 4 of Algorithm 5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
@@ A.2 Reconstruction Function
A.3 Computation of the Bound
To compute the bound of Theorem 3.1, we need three things : the size of the first compression set $\operatorname{\mathbf{i}}$, the size of the second compression set $\operatorname{\mathbf{j}}$ and the message $\mu$. By simply running our CoP2L, we are able to retrieve $|\operatorname{\mathbf{i}}|+|\operatorname{\mathbf{j}}|$, but not the size of each compression set individually. Moreover, we are completely unable to retrieve the messages. Thus, instead of implementing Algorithm 3, we actually implemented the compression function as described in Algorithm 4.
Moreover, we use a block version of P2L, as defined by Algorithm 2 of [23]. We thus add $k$ datapoints at once to the compression set at Line 6 of Algorithm 2 (the used values of $k$ for each experiments is given in Appendix G. Thus, the value of $\mu_{2}^{t}$ (Line 6 of Algorithm 4) becomes the number of iterations of the mP2L algorithm at for task $t$. Knowing that mP2L outputs a compression set $C^{\star}_{t}$, and adds $k$ datapoints to the compression set at each iteration, the number of iterations is $\mu_{2}^{t}=\frac{|C^{\star}_{t}|}{k}$. Finally, when computing the bound, we use the code provided by [65] to invert the $\mathrm{kl}$ divergence.
Appendix B Tighter sample compression $\mathrm{kl}$ bound
In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem 2.1 but proved in a very similar way. To do so, we start by providing the Chernoff test-set bound [66] for losses in $[0,1]$ [67].
Theorem B.1 ([67]).
@@ Appendix B Tighter sample compression $\mathrm{kl}$ bound
$p\leq\mathrm{kl}^{-1}\quantity(\tfrac{1}{n}\textstyle\sum_{i=1}^{n}X_{i},\tfrac{1}{n}\log\tfrac{1}{\delta}).$
We now prove a tighter $\mathrm{kl}$ sample compression bound. As our algorithm CoP2L needs two compression sets, we consider a second compression set $\operatorname{\mathbf{j}}$. We define a conditional probability distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})$ that incorporates the knowledge that a vector $\operatorname{\mathbf{i}}$ was already drawn from $\operatorname{\mathscr{P}}(n)$ and that $\operatorname{\mathbf{i}}\cap\operatorname{\mathbf{j}}=\emptyset$. Thus, we have $\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})\leq 1.$ If the choice of $\operatorname{\mathbf{j}}$ isn’t conditional to the choice of $\operatorname{\mathbf{i}}$, we simply have $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}})$.
Theorem B.2.
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}})\,|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$ , for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
@@
$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
@@
We prove the complement of the expression in Theorem B.2. Denote the upper bound
$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
@@
We use the union bound in Equation 8 and Equation 9. In Equation 10, we use the i.i.d. assumption. In Equation 11 we use the Chernoff Test-set bound of Theorem B.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$. Finally, the last line is derived from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. ∎
Appendix C Proof of Theorem 3.1
@@
Let us choose $t\in[1,T]$. Let us sample $S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}$. We denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$
@@
$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta}).$
Then, we have :
$\operatorname*{\mathbb{P}}_{S^{t}}\quantity(\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$
As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$. Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem B.2. We can then lower bound this probability by $1-\delta$.
@@
$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Biggl{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
@@
As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. Following the work of [39], which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose
@@
Finally, we split the message $\mu$ into two messages $\mu_{1}$ and $\mu_{2}$. The first message is defined using the alphabet $\Sigma_{1}=\{2,\ldots,T\}$. This message indicates the task number where a datapoint is sampled out of the buffer. Thus, the size of $\Sigma_{1}$ is $T-1$ and the probability of a symbol is $\frac{1}{T-1}$. For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
@@
The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,\ldots,n_{t}\}$. We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character. We know that for any $N\geq 0$, $\sum_{k=1}^{N}\zeta(k)\leq 1$. Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\mu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.

### v4

@@ Appendix B Compression and Reconstruction Functions
19:  return $\theta_{T},\{S^{\operatorname{\mathbf{i}}}_{t}\}_{t=1}^{T},\{S^{\operatorname{\mathbf{j}}}_{t}\}_{t=1}^{T},\{\mu_{1,t}\}_{t=1}^{T},\mu_{2}$. {Learned parameters, compression sets and message sets}
B.1 Compression Function
Two challenges prevent CoP2L (as presented by Algorithm 3) from being used as its own compression and reconstruction function, as it is done when working with the original P2L algorithm in a standard (non continual learning) setting. In this subsection, we expose each of these challenges while explaining how the proposed compression algorithm (Algorithm 4) allows to circumvent them.
The first challenge comes from the sampling of the buffer $B^{\star}$ (Lines 6 to 8 of Algorithm 3). When training using CoP2L, datapoints from the buffer can be chosen to be part of the compression set. At the next task, these datapoints may still be available to train the model (if they are not excluded from the buffer by the sampling step). Therefore, the reconstruction function needs to know when to remove the datapoint. The message $\mu_{1}$ returned by Algorithm 4 thus provides the task index after which the datapoint was removed during the execution of CoP2L. As the reconstruction function only requires a message for datapoints who were removed from the buffer after being added to the compression set $S^{\operatorname{\mathbf{i}}}$, a second compression set $S^{\operatorname{\mathbf{j}}}$ is dedicated to datapoints that necessitate a message. The message $\mu_{1}$ given is chosen among $\{2,\ldots,T\}$ for each datapoint belonging to $S^{\operatorname{\mathbf{j}}}$.
The second challenge is the use of the bound as stopping criterion in mP2L. Recall that the reconstruction function does not have access to the whole dataset. Thus, it cannot recover the bound value computed during the initial training phase (see Line 11 of Algorithm˜2) and use it as a stopping criterion. To address this inconvenience, the number of iterations to perform is provided to mP2L as a message $\mu_{2}=(\mu_{2}^{1},\ldots,\mu_{2}^{T})$. That is, for each task $t$, with a buffer $\mathcal{B}$, the message component $\mu_{2}^{t}$ is chosen in $\{1,\ldots,n_{t}+|\mathcal{B}|\}$, giving the number of mP2L’s iterations to perform.
@@ B.1 Compression Function
$\mathscr{M}_{1:T}(\operatorname{\mathbf{j}})\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\textstyle\bigtimes_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (6)
B.2 Reconstruction Function
Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function (Algorithm 4) and for CoP2L (Algorithm 3), the composition of the compression and the reconstruction function must output the exact same predictor as CoP2L. However, it only has access to the compression sets and the messages. The first challenge of the compression function is addressed using $\mu_{1}$ from line 6 to 11 of Algorithm˜5. For each datasets, we verify if a datapoint was previously excluded from the buffer by the sampling function. If so, we remove it from the buffer. The second challenge is addressed in line 4 of Algorithm˜5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
@@ B.2 Reconstruction Function
B.3 Computation of the Bound
To compute the bound of Theorem˜3.1, we need three things : the size of the first compression set $\operatorname{\mathbf{i}}$, the size of the second compression set $\operatorname{\mathbf{j}}$ and the message $\mu$. By simply running our CoP2L, we are able to retrieve $|\operatorname{\mathbf{i}}|+|\operatorname{\mathbf{j}}|$, but not the size of each compression set individually. Moreover, we are completely unable to retrieve the messages. Thus, instead of implementing Algorithm˜3, we actually implemented the compression function as described in Algorithm˜4.
Moreover, we use a block version of P2L, as defined by Algorithm 2 of Paccagnan et al. [2024]. We thus add $k$ datapoints at once to the compression set at Line 6 of Algorithm 2 (the used values of $k$ for each experiments is given in Section˜E.2. Thus, the value of $\mu_{2}^{t}$ (Line 6 of Algorithm˜4) becomes the number of iterations of the mP2L algorithm at for task $t$. Knowing that mP2L outputs a compression set $C^{\star}_{t}$, and adds $k$ datapoints to the compression set at each iteration, the number of iterations is $\mu_{2}^{t}=\frac{|C^{\star}_{t}|}{k}$. Finally, when computing the bound, we use the code provided by Viallard et al. [2021] to invert the $\mathrm{kl}$ divergence.
Appendix C Tighter sample compression $\mathrm{kl}$ bound
In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem˜2.1 but proved in a very similar way. To do so, we start by providing the Chernoff test-set bound for losses in $[0,1]$ [Langford, 2005, Foong et al., 2022].
Theorem C.1 (Foong et al., 2022).
@@ Appendix C Tighter sample compression $\mathrm{kl}$ bound
[…]
We now prove a tighter $\mathrm{kl}$ sample compression bound. As our algorithm CoP2L needs two compression sets, we consider a second compression set $\operatorname{\mathbf{j}}$. To use these two compression sets, we need to first redefine the reconstruction function
$\operatorname{\mathscr{R}}:\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{k\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{k}\times\boldsymbol{\cup}_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{j}})\to\Theta.$
We define a conditional probability distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})$ that incorporates the knowledge that a vector $\operatorname{\mathbf{i}}$ was already drawn from $\operatorname{\mathscr{P}}(n)$ and that $\operatorname{\mathbf{i}}\cap\operatorname{\mathbf{j}}=\emptyset$. Thus, we have $\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})\leq 1.$ If the choice of $\operatorname{\mathbf{j}}$ isn’t conditional to the choice of $\operatorname{\mathbf{i}}$, we simply have $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}})$.
Theorem C.2.
[…]
@@
$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\ln\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
@@
Let us prove the complement of the expression in Theorem˜C.2. Denote the upper bound
$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\ =\ \mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\ln\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})\,.$
We have
@@
Equations˜7 and 8 use the union bound, Equation˜9 uses the i.i.d. assumption, and Equation˜10 uses the Chernoff Test-set bound of Theorem˜C.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$. Finally, the last inequality is obtained from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. ∎
Appendix D Proof of Theorem 3.1
@@
Let us choose $t\in[1,T]$. Let us sample $S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}$.
@@
$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)}),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\ln\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta}).$
Then, we have
$\operatorname*{\mathbb{P}}_{S^{t}}\quantity(\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)).$
As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$. Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem˜C.2. We can then lower bound this probability by $1-\delta$.
@@
$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Biggl(\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr)$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg(\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr)$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg(\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr)$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg(\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr)$
@@
As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. Following the work of Marchand et al. [2003], which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose
@@
Finally, we split the message $\mu$ into two messages $\mu_{1}$ and $\mu_{2}$. The first message is defined using the alphabet $\Sigma_{1}=\{2,\ldots,T\}$. This message indicates the task number where a datapoint is sampled out of the buffer. Thus, the size of $\Sigma_{1}$ is T1T-1 and the probability of a symbol is $\frac{1}{T-1}$. For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
@@
The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,\ldots,n_{t}+|{\mathcal{B}}|\}$. We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character. We know that for any $N\geq 0$, $\sum_{k=1}^{N}\zeta(k)\leq 1$. Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\mu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.